Here is my question:
Arturo is driving at a speed of 50mph. At 2:00 he sees the following sign.
Applebee 11 mi
Beauville 41 mi
Caanan 62mi
Denton 98mi
Arturo continues at the same speed. At 2:30 how far from Caanan will he be?
A. 12mi
B. 16mi
C. 25mi.
D. 26mi.
E 37mi.

I plugged in the formula for d=rt and got fifteen miles, then subtracted this from the distance to Cannan and got 47mi. The answer is 37mi. E
Can you please let me know where I went wrong here?
Thank you.

Answer by bucky(2189)

At 50 mph, in 30 minutes (1/2 hour) you would go 25 miles, not the 15 miles you got.
.
d = r*t
.
d = 50*(1/2) = 50/2 = 25 miles
.
Since it was 62 miles to Caanan when he started and then he drove 25 miles toward Caanan,
the remaining distance he has to go is 62 - 25 = 37 miles. E is the correct answer.
.
Hope this helps you.